OpenStudy (anonymous):

Of 50 students selected at random at a school, 15 said they preferred their hamburgers plain. How many of the 1500 students enrolled in the school would you expect to prefer their hamburgers plain?

OpenStudy (anonymous):

15/50 = x/1500 3/10 = x/1500 x = 450